# The inverse initial data problem for anisotropic Navier–Stokes equations via Legendre time reduction method

### v2

@@ Abstract
We consider the inverse initial data problem for the compressible anisotropic Navier-Stokes equations, where the goal is to reconstruct the initial velocity field from lateral boundary observations. This problem arises in applications where direct measurements of internal fluid states are unavailable. We introduce a novel computational framework based on Legendre time reduction, which projects the velocity field onto an exponentially weighted Legendre basis in time. This transformation reduces the original time-dependent inverse problem to a coupled, time-independent elliptic system. The resulting reduced model is solved iteratively using a Picard iteration and a stabilized least-squares formulation under noisy boundary data. Numerical experiments in two dimensions confirm that the method accurately and robustly reconstructs initial velocity fields, even in the presence of significant measurement noise and complex anisotropic structures. This approach offers a flexible and computationally tractable alternative for inverse modeling in fluid dynamics with anisotropic media.
@@
[label2]organization=Department of Mathematics, NC State University,city=Raleigh, postcode=27695, state=NC, country=USA
@@ 1 Introduction
Let $d\geq 2$ denote the spatial dimension, and let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with smooth boundary $\partial\Omega$. We consider a fourth-order tensor field $\bm{\mu}:\overline{\Omega}\to\mathbb{R}^{d\times d\times d\times d}$ representing the viscosity property of the medium, which satisfies the coercivity: There exists a constant $\Lambda>0$ such that for all matrices $\bm{\xi}=(\xi_{ij})_{i,j=1}^{d}\in\mathbb{R}^{d\times d}$, we have
@@
Let $T>0$ be a final time. Consider a fluid velocity field $\mathbf{u}:\Omega\times(0,T)\to\mathbb{R}^{d}$ governed by the following initial value problem for the compressible anisotropic Navier–Stokes equations:
$\begin{cases}\displaystyle\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\nabla p+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}\right),&({\bf x},t)\in\Omega\times(0,T),\\ {\bf u}({\bf x},t)=0,&({\bf x},t)\in\partial\Omega\times(0,T),\\ \mathbf{u}({\bf x},0)=\mathbf{U}_{0}({\bf x}),&{\bf x}\in\Omega,\end{cases}$ (1.1)
where $\mathbf{U}_{0}:\Omega\to\mathbb{R}^{d}$ represents the initial velocity field and $p:\Omega\times(0,T)\to\mathbb{R}$ is the fluid pressure. Our focus is on the following inverse problem:
@@
$\mathbf{f}({\bf x},t)=\partial_{\nu}\mathbf{u}({\bf x},t),\quad\mbox{for all }({\bf x},t)\in\partial\Omega\times(0,T),$ (1.2)
reconstruct the initial condition $\mathbf{U}_{0}$.
@@
Our strategy for solving the inverse problem associated with the Navier–Stokes equations stated in Problem 1 is grounded in the use of the Legendre time-reduction technique. Specifically, we approximate the velocity field $\mathbf{u}$ in the time variable by truncating its Fourier series with respect to an exponentially weighted Legendre basis of the weighted space $L_{e^{-2t}}^{2}(0,T)$. This basis, recently introduced in [36] for inverse initial data problems in elasticity, and further utilized in [23] for Maxwell’s system, offers notable advantages over classical bases such as the trigonometric system and the Legendre polynomial without the exponential weight. In particular, the exponential-weighted Legendre polynomials possess two desirable properties: non-vanishing derivatives and favorable convergence behavior of the truncated series applied to time derivatives of the velocity field. By applying this time-reduction procedure, we reformulate the original time-dependent inverse problem into a sequence of time-independent systems, where the principal operator takes the form ${\rm div}(\bm{\mu}:\nabla\cdot)$. We refer to this reformulated system as the time-dimensional reduction model. Leveraging this structure, we then employ a Picard iteration to solve this model. The obtained solution directly yields an approximate solution to the original inverse problem, thereby enabling a globally convergent reconstruction scheme.
@@ 3 The time-dimensional reduction method
The inverse problem stated in Problem 1 can be reformulated as the recovery of a vector field $\mathbf{u}$ satisfying
$\begin{cases}\partial_{t}\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\nabla p+{\rm div}(\bm{\mu}:\nabla\mathbf{u})&\text{in }\Omega\times(0,T),\\ \mathbf{u}({\bf x},t)=\textbf{0}&\text{on }\partial\Omega\times(0,T),\\ \partial_{\nu}\mathbf{u}({\bf x},t)=\mathbf{f}({\bf x},t)&\text{on }\partial\Omega\times(0,T),\end{cases}$ (3.1)
given the data $\mathbf{f}$ as in (1.2).
@@
The existence of a solution to (3.1) is guaranteed, as $\mathbf{u}$ corresponds to the restriction of the solution to the forward problem (1.1) on $\Omega\times(0,T)$. Although a rigorous proof of uniqueness for the original inverse problem is beyond the scope of this work, which is primarily concerned with computational methodologies, we emphasize that this question remains of significant theoretical interest.
We now derive the time-dimensional reduction model. Let $N\geq 1$ be a fixed integer, and consider the projection operator $\mathbb{P}^{N}$. Applying $\mathbb{P}^{N}$ to both sides of the differential equation in (3.1), we obtain:
$\mathbb{P}^{N}[\partial_{t}\mathbf{u}^{*}]+\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]=-\mathbb{P}^{N}[\nabla p]+{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}])\quad\mbox{in }\Omega\times(0,T).$ (3.2)
Adding and subtracting $\partial_{t}\mathbb{P}^{N}[{\bf u}]+(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]$ on the left-hand side of (3.2) and then rearranging the resulting equation yields
$\partial_{t}\mathbb{P}^{N}[\mathbf{u}]-{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}])+(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]+\mathbb{P}^{N}[\nabla p]\\ =-\left(\mathbb{P}^{N}[\partial_{t}\mathbf{u}]-\partial_{t}\mathbb{P}^{N}[\mathbf{u}]\right)-\left(\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]-(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]\right)$ (3.3)
in $\Omega\times(0,T)$. We have the theorem
@@
Assume the solution $\bf u$} to (3.1) belongs to the class $H^{3}((0,T),H^{2}(\Omega)^{d})$. Then,
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[\partial_{t}\mathbf{u}]-\partial_{t}\mathbb{P}^{N}[\mathbf{u}]\right\|_{L^{2}((0,T);H^{2}(\Omega)^{d})}=0$ (3.4)
@@
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]-(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]\right\|_{H^{2}(\Omega)^{d}}=0.$ (3.5)
@@
We first establish the convergence of the series $\sum_{n=0}^{\infty}\left\langle{\bf u}({\bf x},\cdot),\Psi_{n}\right\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ in $L^{2}(\Omega)^{d}.$ It follows from (2.1) for the vector $\bf u$} and for $k=3$ that for each ${\bf x}\in\Omega$
$\sum_{n=0}^{\infty}n^{6}\left|\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)^{d}}\right|^{2}\leq C\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)^{d}}^{2}.$ (3.6)
Since each term of the series in the left-hand side of (3.6) is non-negative, we have for all $n\geq 1$,
$n^{6}\left|\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)^{d}}\right|^{2}\leq C\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)^{d}}^{2}.$ (3.7)
Integrating both sides of (3.7) gives
@@
$n^{6}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}\leq C\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}.$ (3.8)
Therefore, multiplying $e^{-2t}|\Psi_{n}^{\prime}(t)|^{2}$ to both sides of (3.8) and integrating the resulting inequality, we obtain
$\int_{0}^{T}n^{6}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt$ $\leq C\int_{0}^{T}e^{-2t}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}|\Psi^{\prime}(t)|^{2}d{\bf x}dt$
$\leq C\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}\int_{0}^{T}e^{-2t}|\Psi^{\prime}_{n}(t)|^{2}dt$
$\leq Cn^{3}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}.$
Here, we have used (2.2) at the last estimate. Hence,
$\int_{0}^{T}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt\leq Cn^{-3}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}$ (3.9)
for $n\geq 1$. Here, we avoid the case $n=0$ because of the appearance of $n^{-3}$ on the right-hand side of (3.9). It follows from (3.9) that
$\sum_{n=1}^{\infty}\Big(\int_{0}^{T}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt\Big)^{1/2}\\ \leq C\sum_{n=1}^{\infty}n^{-3/2}\Big(\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}\Big)^{1/2}<\infty.$ (3.10)
It follows from (3.10) that the series $\sum_{n=0}^{\infty}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ Define the limit function
${\bf w}({\bf x},t):=\lim_{N\to\infty}\sum_{n=0}^{N}\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\,\Psi_{n}^{\prime}(t).$
We show that ${\bf w}={\bf u}_{t}$. Let $\varphi\in C_{c}^{\infty}(\Omega\times(0,T))^{d}$ be an arbitrary test function. We compute:
$\int_{\Omega}\int_{0}^{T}{\bf w}({\bf x},t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$ $=\int_{\Omega}\int_{0}^{T}\lim_{N\to\infty}\left(\sum_{n=0}^{N}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)\right)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$
$=\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\int_{0}^{T}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\,\Psi_{n}^{\prime}(t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t.$
This interchange of the limit and the integral above is justified by the convergence with respect to the norm of $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$ of the partial sums to $\bf w$}, while $\varphi$ is compactly supported and bounded. Next, we integrate by parts in time:
$\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\int_{0}^{T}\langle{\bf u}({\bf x},\cdot),\Psi_{n}$ $\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$
$=\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\cdot\left(\int_{0}^{T}\Psi_{n}^{\prime}(t)\varphi({\bf x},t){\rm d}t\right){\rm d}{\bf x}$
$=-\int_{\Omega}\lim_{N\to\infty}\sum_{n=0}^{N}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\cdot\left(\int_{0}^{T}\Psi_{n}(t)\varphi_{t}({\bf x},t){\rm d}t\right){\rm d}{\bf x}$
$=-\int_{\Omega}\int_{0}^{T}{\bf u}({\bf x},t)\cdot\varphi_{t}({\bf x},t)\,{\rm d}{\bf x}{\rm d}t,$
where we used completeness of the $\{\Psi_{n}\}$ basis in $L^{2}_{e^{-2t}}(0,T)$, convergence of the series, and the dominated convergence theorem again. Hence, ${\bf w}={\bf u}_{t}$ in the distributional sense, and since both of these vector-valued functions lie in $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$, the equality holds almost everywhere.
We can repeat this arguments for $\partial_{x_{i}}{\bf u}({\bf x},t)$ and $\partial_{x_{i}x_{j}}{\bf u}({\bf x},t)$, $1\leq i,j\leq d$ to obtain the convergences of $\sum_{n=0}^{\infty}\langle\partial_{x_{i}}{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ and $\sum_{n=0}^{\infty}\langle\partial_{x_{i}x_{j}}{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ to $\partial_{x_{i}}{\bf w}$ and $\partial_{x_{i}x_{j}}{\bf w}$ respectively in $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$. We obtain (3.4).
Proving (3.5) is straightforward because the projection operator $\mathbb{P}^{N}$ acts only on the temporal variable $t$ and does not affect the spatial variable $\bf x$}. ∎
Remark 4 (Regularity in Theorem 3.5 and observed performance).
Theorem 3.5 assumes high regularity (for example, $\mathbf{u}\in H^{3}(0,T;H^{2}(\Omega)^{d})$). These assumptions are used to carry out the analysis and obtain clear statements with explicit bounds.
In computations, the algorithm performs well even when the initial condition $\mathbf{u}_{0}$ is only in $L^{2}(\Omega)^{d}$ and may be discontinuous, so the hypotheses of Theorem 3.5 do not hold, see Section 5 for our out-of-expectation numerical results. This indicates that the method’s practical scope is broader than current theory covers.
Due to Theorem 3.5 and the limits (3.4) and (3.5), equation (3.3) can be approximated by the reduced system
$\partial_{t}\mathbb{P}^{N}[\mathbf{u}]+(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]=-\mathbb{P}^{N}[\nabla p]+{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}]),$ (3.11)
which serves as the foundation of our time-dimensional reduction model. For each $n\geq 0$, let
$\mathbf{u}_{n}({\bf x})=\int_{0}^{T}e^{-2t}\mathbf{u}({\bf x},t)\Psi_{n}(t)\,dt,$ (3.12)
be the $n$-th Fourier mode of $\mathbf{u}$ with respect to the basis $\{\Psi_{n}\}_{n\geq 0}$.

### v4

@@ Abstract
We consider an inverse initial-data problem for the compressible anisotropic Navier–Stokes equations, in which the goal is to reconstruct the initial velocity field from noisy lateral boundary observations. In the formulation studied here, the density, pressure, anisotropic viscosity tensor, and body force are assumed known, while the initial velocity is the quantity to be recovered. We introduce a new computational framework based on Legendre time-dimensional reduction, in which the velocity field is projected onto an exponentially weighted Legendre basis in time. This transformation reduces the original time-dependent inverse problem to a coupled system of time-independent elliptic equations for the Fourier coefficients of the velocity field. The resulting reduced model is solved using a combination of quasi-reversibility and a damped Picard iteration. Numerical experiments in two dimensions show that the proposed method accurately and robustly reconstructs initial velocity fields, even in the presence of significant measurement noise, geometrically complex structures, and anisotropic effects. The method provides a flexible and computationally tractable approach for inverse fluid problems in anisotropic media.
@@
[label2]organization=Department of Mathematics and Statistics, California State University, Long Beach,city=Long Beach, postcode=90840, state=CA, country=USA
@@ 1 Introduction
Let $d\geq 2$ denote the spatial dimension, and let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with smooth boundary $\partial\Omega$. We consider a fourth-order tensor field $\bm{\mu}:\overline{\Omega}\to\mathbb{R}^{d\times d\times d\times d}$ representing the viscosity property of the medium, which satisfies the coercivity condition: there exists a constant $\Lambda>0$ such that for all matrices $\bm{\xi}=(\xi_{ij})_{i,j=1}^{d}\in\mathbb{R}^{d\times d}$, we have
@@
Let $T>0$ be a final time. We consider a fluid velocity field ${\bf u}:\Omega\times(0,T)\to\mathbb{R}^{d}$ governed by the following initial–boundary value problem for the compressible anisotropic Navier–Stokes system:
$\begin{cases}\displaystyle\rho\left(\frac{\partial{\bf u}}{\partial t}+({\bf u}\cdot\nabla){\bf u}\right)=-\nabla p+{\rm div}\left(\bm{\mu}:\nabla{\bf u}\right)+\rho({\bf x},t){\bf F}({\bf x},t),&({\bf x},t)\in\Omega\times(0,T),\\[3.99994pt] \rho_{t}+{\rm div}(\rho{\bf u})=0,&({\bf x},t)\in\Omega\times(0,T),\\[3.99994pt] p=c^{2}\rho\qquad\text{or}\qquad p=K\rho^{\gamma},&({\bf x},t)\in\Omega\times(0,T),\\[3.99994pt] {\bf u}({\bf x},t)=0,&({\bf x},t)\in\partial\Omega\times(0,T),\\[3.99994pt] {\bf u}({\bf x},0)={\bf u}^{0}({\bf x}),&{\bf x}\in\Omega,\\[3.99994pt] \rho({\bf x},0)=\rho^{0}({\bf x}),&{\bf x}\in\Omega.\end{cases}$ (1.1)
Here, ${\bf u}^{0}:\Omega\to\mathbb{R}^{d}$ denotes the initial velocity field, $\rho:\Omega\times(0,T)\to\mathbb{R}$ is the density, $\rho^{0}({\bf x})$ is a given positive initial density, $p:\Omega\times(0,T)\to\mathbb{R}$ is the pressure, and ${\bf F}:\Omega\times(0,T)\to\mathbb{R}^{d}$ is the body force. The constants $c>0$, $K>0$, and $\gamma>1$ are given, where $p=c^{2}\rho$ corresponds to the isothermal case and $p=K\rho^{\gamma}$ corresponds to the polytropic law.
Our focus is on the following inverse problem:
@@
$\mathbf{f}({\bf x},t)=\partial_{\nu}\mathbf{u}({\bf x},t)\quad\mbox{for all }({\bf x},t)\in\partial\Omega\times(0,T),$ (1.2)
assume that the density $\rho({\bf x},t)$, the pressure $p({\bf x},t)$, the anisotropic viscosity tensor $\bm{\mu}$, and the body force ${\bf F}({\bf x},t)$ are known. Reconstruct the initial condition ${\bf u}^{0}({\bf x})$ in $\Omega$.
To avoid ambiguity in the formulation of the inverse problem, we now clarify which quantities are intended to be reconstructed and which are treated as known quantities. Since the observation (1.2) provides only boundary information, attempting to reconstruct multiple independent space-time fields from this data would be severely underdetermined without supplementary measurements or additional structural assumptions. We emphasize that the simultaneous recovery of quantities such as $\rho({\bf x},t)$, $p({\bf x},t)$, and ${\bf F}({\bf x},t)$ would be mathematically significant and practically very interesting. However, in the present setting, the observation is given only on a (d1)(d-1)-dimensional spatial manifold over time, making the total dimension (d1)+1=d(d-1)+1=d, whereas these unknown fields are defined in the full $d$-dimensional space-time cylinder $\Omega\times(0,T)$, whose dimension is $d+1$. Therefore, the available data do not contain enough information to identify all these quantities simultaneously. Accordingly, we formulate a targeted inverse problem in which the initial condition ${\bf u}^{0}$ is the sole reconstruction target, while $\rho({\bf x},t)$, $p({\bf x},t)$, $\bm{\mu}$, and ${\bf F}({\bf x},t)$ are assumed known.
@@
Unlike the standard PDE-constrained optimization framework based on state, adjoint, and optimality systems, our approach follows a direct reconstruction route. Our strategy for solving the inverse problem associated with the Navier–Stokes equations stated in Problem 1 is grounded in the use of the Legendre time-reduction technique. Specifically, we approximate the velocity field $\mathbf{u}$ in the time variable by truncating its Fourier series with respect to an exponentially weighted Legendre basis of the weighted space $L_{e^{-2t}}^{2}(0,T)$. This basis, recently introduced in [37] for inverse initial data problems in elasticity, and further utilized in [23] for Maxwell’s system, offers notable advantages over classical bases such as the trigonometric system and the Legendre polynomial basis without the exponential weight. In particular, the exponentially weighted Legendre basis possesses two desirable properties: nonvanishing derivatives and favorable convergence behavior of the truncated series applied to time derivatives of the velocity field. By applying this time-reduction procedure, we reformulate the original time-dependent inverse problem into a coupled system of time-independent equations, where the principal operator takes the form ${\rm div}(\bm{\mu}:\nabla\cdot)$. We refer to this reformulated system as the time-dimensional reduction model. Leveraging this structure, we then employ quasi-reversibility together with a damped Picard iteration to solve the reduced model. The obtained solution directly yields an approximate solution to the original inverse problem, thereby providing a practical iterative reconstruction scheme.
@@ 3 The time-dimensional reduction method
The inverse problem stated in Problem 1 can be reformulated as the recovery of a vector field $\bf u$} satisfying
$\begin{cases}\displaystyle\rho({\bf x},t)\left(\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}\right)=-\nabla p({\bf x},t)+{\rm div}(\bm{\mu}:\nabla{\bf u})+\rho({\bf x},t){\bf F}({\bf x},t),&\text{in }\Omega\times(0,T),\\[3.99994pt] {\bf u}({\bf x},t)=\mathbf{0},&\text{on }\partial\Omega\times(0,T),\\[3.99994pt] \partial_{\nu}{\bf u}({\bf x},t)=\mathbf{f}({\bf x},t),&\text{on }\partial\Omega\times(0,T),\end{cases}$ (3.1)
where $\rho({\bf x},t)$, $p({\bf x},t)$, $\bm{\mu}$, and ${\bf F}({\bf x},t)$ are known, while the initial condition ${\bf u}^{0}$ is the unknown to be reconstructed.
@@
The existence of a solution to (3.1) is justified by the fact that $\bf u$} is the restriction of the solution to the forward problem (1.1) on $\Omega\times(0,T)$. A rigorous analysis of uniqueness for the original inverse problem is beyond the scope of this paper, whose main focus is on the computational methodology. Nevertheless, this question remains of independent theoretical interest.
Remark 4.
The continuity equation and any closure relation between $p$ and $\rho$ (such as $p=c^{2}\rho$ or $p=K\rho^{\gamma}$) belong to the forward compressible model. However, they are not explicitly needed in the reconstruction step considered in this paper. Indeed, in our targeted inverse problem, the density $\rho({\bf x},t)$ and the pressure $p({\bf x},t)$ are assumed known, together with $\bm{\mu}$ and ${\bf F}({\bf x},t)$. Therefore, the momentum equation alone forms a closed system of $d$ equations for the $d$ components of the unknown velocity field $\bf u$}. For this reason, the time-dimensional reduction procedure is applied only to the momentum equation.
We now derive the time-dimensional reduction model. Let $N\geq 1$ be a fixed integer, and consider the projection operator $\mathbb{P}^{N}$. Applying $\mathbb{P}^{N}$ to both sides of the differential equation in (3.1), we obtain
$\mathbb{P}^{N}\!\left[\rho({\bf x},t)\partial_{t}{\bf u}\right]+\mathbb{P}^{N}\!\left[\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}\right]=-\mathbb{P}^{N}[\nabla p({\bf x},t)]+{\rm div}\left(\bm{\mu}:\nabla\mathbb{P}^{N}[{\bf u}]\right)+\mathbb{P}^{N}\!\left[\rho({\bf x},t){\bf F}({\bf x},t)\right]$ (3.2)
in $\Omega\times(0,T)$, where we used the fact that $\bm{\mu}$ is known and independent of $t$. In (3.2), we have used the identity
$\mathbb{P}^{N}[{\rm div}(\bm{\mu}:\nabla{\bf u})]={\rm div}\bigl(\bm{\mu}:\nabla\mathbb{P}^{N}[{\bf u}]\bigr).$
This is true because $\mathbb{P}^{N}$ acts only on the temporal variable and ${\rm div}(\bm{\mu}:\nabla,\cdot,)$ is a spatial operator with time-independent coefficient tensor $\bm{\mu}$.
Adding and subtracting $\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]+\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]$ on the left-hand side of (3.2), and then rearranging the resulting equation, we obtain
$\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]-{\rm div}\left(\bm{\mu}:\nabla\mathbb{P}^{N}[{\bf u}]\right)+\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\\ +\mathbb{P}^{N}[\nabla p({\bf x},t)]-\mathbb{P}^{N}[\rho({\bf x},t){\bf F}({\bf x},t)]=-\Bigl(\mathbb{P}^{N}[\rho({\bf x},t)\partial_{t}{\bf u}]-\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]\Bigr)\\ -\Bigl(\mathbb{P}^{N}[\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}]-\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\Bigr)$ (3.3)
in $\Omega\times(0,T)$.
We have the theorem
@@
Assume that the solution $\bf u$} to (3.1) satisfies ${\bf u}\in H^{3}\bigl((0,T);H^{s}(\Omega)^{d}\bigr)$ for some integer $s>\frac{d}{2}+1$. Then
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[\partial_{t}{\bf u}]-\partial_{t}\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}=0,$ (3.4)
@@
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[({\bf u}\cdot\nabla){\bf u}]-(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}=0.$ (3.5)
@@
Since the weighted norm on $L^{2}_{e^{-2t}}(0,T)$ is equivalent to the standard $L^{2}(0,T)$ norm on the finite interval $(0,T)$, we will work with the standard time norms below.
We first prove (3.4). By the approximation property of the projection operator $\mathbb{P}^{N}$ associated with the weighted Legendre basis, we have
$\mathbb{P}^{N}[{\bf u}]\to{\bf u}\quad\text{in }H^{1}\bigl((0,T);H^{s}(\Omega)^{d}\bigr),$
and
$\mathbb{P}^{N}[\partial_{t}{\bf u}]\to\partial_{t}{\bf u}\quad\text{in }L^{2}\bigl((0,T);H^{s}(\Omega)^{d}\bigr).$
Therefore,
$\left\|\mathbb{P}^{N}[\partial_{t}{\bf u}]-\partial_{t}\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}$ $\leq\left\|\mathbb{P}^{N}[\partial_{t}{\bf u}]-\partial_{t}{\bf u}\right\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}$
$\quad+\left\|\partial_{t}{\bf u}-\partial_{t}\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}.$
Both terms on the right-hand side converge to zero as $N\to\infty$, which proves (3.4).
We next prove (3.5). Define the nonlinear mapping
$\mathcal{N}({\bf v}):=({\bf v}\cdot\nabla){\bf v}.$
Since $s>\frac{d}{2}$, the product of two functions in $H^{s}(\Omega)$ still belongs to $H^{s}(\Omega)$. Moreover, since $s>\frac{d}{2}+1$, the mapping
$\mathcal{N}:H^{s}(\Omega)^{d}\to H^{s-1}(\Omega)^{d}$
is locally Lipschitz. In particular, there exists a constant $C>0$ such that for all ${\bf v},{\bf w}\in H^{s}(\Omega)^{d}$,
$\|\mathcal{N}({\bf v})-\mathcal{N}({\bf w})\|_{H^{s-1}(\Omega)^{d}}\leq C\bigl(\|{\bf v}\|_{H^{s}(\Omega)^{d}}+\|{\bf w}\|_{H^{s}(\Omega)^{d}}\bigr)\|{\bf v}-{\bf w}\|_{H^{s}(\Omega)^{d}}.$ (3.6)
Now write
$\left\|\mathbb{P}^{N}[({\bf u}\cdot\nabla){\bf u}]-(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}$
$\qquad\leq\left\|\mathbb{P}^{N}[({\bf u}\cdot\nabla){\bf u}]-({\bf u}\cdot\nabla){\bf u}\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}$
@@
$\qquad\quad+\left\|({\bf u}\cdot\nabla){\bf u}-(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}.$
The first term tends to zero because
$({\bf u}\cdot\nabla){\bf u}\in L^{2}\bigl((0,T);H^{s-1}(\Omega)^{d}\bigr)$
and $\mathbb{P}^{N}$ converges to the identity in this space.
For the second term, applying (3.6) pointwise in time and then integrating over $(0,T)$, we obtain
$\left\|({\bf u}\cdot\nabla){\bf u}-(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}$
$\qquad\leq C\Bigl(\|{\bf u}\|_{L^{\infty}((0,T);H^{s}(\Omega)^{d})}+\|\mathbb{P}^{N}[{\bf u}]\|_{L^{\infty}((0,T);H^{s}(\Omega)^{d})}\Bigr)\|{\bf u}-\mathbb{P}^{N}[{\bf u}]\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}.$
Since
$\mathbb{P}^{N}[{\bf u}]\to{\bf u}\quad\text{in }H^{1}\bigl((0,T);H^{s}(\Omega)^{d}\bigr),$
the Sobolev embedding
$H^{1}\bigl((0,T);H^{s}(\Omega)^{d}\bigr)\hookrightarrow C\bigl([0,T];H^{s}(\Omega)^{d}\bigr)$
implies that $\{\mathbb{P}^{N}[{\bf u}]\}_{N\geq 1}$ is uniformly bounded in $L^{\infty}((0,T);H^{s}(\Omega)^{d})$, and
$\|{\bf u}-\mathbb{P}^{N}[{\bf u}]\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}\to 0.$
Hence, the second term also converges to zero. This proves (3.5). ∎